### v1

@@ Abstract
This article investigates the integration of quasi-Monte Carlo (QMC) methods using the Adaptive Multiple Importance Sampling (AMIS). Traditional Importance Sampling (IS) often suffers from poor performance since it heavily relies on the choice of the proposal distributions. The AMIS and the Modified version of AMIS (MAMIS) address this by iteratively refining proposal distributions and reusing all past samples through a recycling strategy. We introduce the RQMC methods into the MAMIS, achieving higher convergence rates compared to the Monte Carlo (MC) methods. Our main contributions include a detailed convergence analysis of the MAMIS estimator under randomized QMC (RQMC) sampling. Specifically, we establish the $L^{q}$ $(q\geq 2)$ error bound for the RQMC-based estimator using a smoothed projection method, which enables us to apply the Hölder’s inequality in the error analysis of the RQMC-based MAMIS estimator. As a result, we prove that the root mean square error of the RQMC-based MAMIS estimator converges at a rate of $\mathcal{O}(\bar{N}_{T}^{-1+\epsilon})$, where $\bar{N}_{T}$ is the average number of samples used in each step over $T$ iterations, and $\epsilon>0$ is arbitrarily small. Numerical experiments validate the effectiveness of our method, including mixtures of Gaussians, a banana-shaped model, and Bayesian Logistic regression.
@@ 2.1 The modified version of AMIS
Utilizing the samples $X_{i}^{T}$ with weights $\omega_{i}^{T}$ obtained at the final stage $T$, the AIS estimator is formulated as
@@
Marin et al. [15] also emphasized this recycling strategy, but instead of reweighting the samples at each stage $t$ to update the parameters, a simpler and more analytically tractable method was suggested. The resulting Modified version of AMIS algorithm is exposed in Algorithm 2. We use the assumption in [15], that is, whatever the criterion we believe in (Kullback-Leibler divergence, variance, or others), the parameter $\theta^{\ast}$ which we desire $\widehat{\theta}_{t}$ to approximate, can be written as
@@ 3 Convergence rate of the RQMC-based MAMIS estimator
$\Big{(}\mathbb{E}\Big{|}\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I\Big{|}^{2}\Big{)}^{\frac{1}{2}}\leq C(M,B,k,d,D,L,\eta,\epsilon)\bar{N}_{T}^{-1+\epsilon},$
where the constant $C$ depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@
$\left(\mathbb{E}|\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I|^{2}\right)^{\frac{1}{2}}\leq C(M,B,k,d,D,L,\eta,\epsilon)N^{-1+\epsilon},$
where the constant $C$ depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@
Unfortunately, this corollary does not include the relationship between error and the number of iteration $T$. This is because $T$ actually appears in the constant term. In order to estimate the sum of errors at each stage, we use the same $G_{e}(M,B,k)$ to encompass all ${\pi\psi}/{q(\cdot,\theta^{\ast})}\circ T_{\theta},{\pi h_{j}}/{q(\cdot,\theta^{\ast})}\circ T_{\theta}$, which eliminated the influence of $T$ in the constant term. Moreover, Theorem 3.1 is not entirely devoid of the impact of $T$. When we use a strictly increasing sample size $\{N_{t}\}$, such as $N_{t}=Nt$, we have $\bar{N}_{T}=\frac{N(T+1)}{2}$. Applying Theorem 3.1 yields an error order as $\mathcal{O}((NT)^{-1+\epsilon})$. {Remark} The convergence order in Theorem 3.1 still holds for the case where the integrand $\psi$ is a finite-dimensional vector-valued function. This can be shown by applying Theorem 3.1 to each component function of $\psi$ and noting that the square of the vector norm is equal to the sum of the squares of its components. In this case, the constant term in Theorem 3.1 will also be related to the dimension of the vector-valued function.
@@ 3.1 Convergence of the parameters $\widehat{\theta}_{t}$
$\left(\mathbb{E}{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\right)^{\frac{1}{q}}\leq C(M,B,k,d,q)N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(4qd\log N_{t})^{\frac{k}{2}}\}}.$
Thus we can obtain
(14) $\left(\mathbb{E}{\lVert\widehat{\theta}_{t+1}-\theta^{\ast}\rVert}^{q}\right)^{\frac{1}{q}}$ $\leq C(M,B,k,d,D,q)N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(4qd\log N_{t})^{\frac{k}{2}}\}}.$
@@
From Eq. 13, the conditional expectation $\mathbb{E}\left[{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\Big{|}\widehat{\theta_{t}}\right]$ is equal to
$\mathbb{E}\left[\Big{|}\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}h_{\widehat{\theta}_{t}}^{(j)}\circ F_{\widehat{\theta}_{t}}(u_{i}^{t})-\int_{[0,1)^{d}}h_{\widehat{\theta}_{t}}^{(j)}\circ F_{\widehat{\theta}_{t}}(\bm{u})\mathrm{d}\bm{u}\Big{|}^{q}\bigg{|}\widehat{\theta}_{t}\right]$
@@
$\leq$ $C(M,B,k,d,q)N_{t}^{-q}(\log N_{t})^{q(\frac{3d}{2}-1)}\exp{\{qM(4qd\log N_{t})^{\frac{k}{2}}\}}$
for any $\theta\in\Theta$, where Eq. 15 employs the independence between the RQMC point set $\{u_{i}^{t}\}_{i=1}^{N_{t}}$ and $\widehat{\theta}_{t}$. Thus we obtain
$\mathbb{E}\left[{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\Big{|}\widehat{\theta_{t}}\right]\leq C(M,B,k,d,q)N_{t}^{-q}(\log N_{t})^{q(\frac{3d}{2}-1)}\exp{\{qM(4qd\log N_{t})^{\frac{k}{2}}\}}.\$
@@ 3.2 On the property of operator $I_{\psi}^{\ast}$
$\Big{(}\mathbb{E}\big{|}I_{\psi}^{\ast}(\widehat{\theta}_{t})-I_{\psi}^{\ast}(\theta^{\ast})\big{|}^{2}\Big{)}^{\frac{1}{2}}\leq C(M,B,k,d,D,L)N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}.$
@@ 3.3 Convergence of the auxiliary estimator
$\left(\mathbb{E}|\widehat{\pi}_{t}^{\ast}(\psi)-I_{\psi}^{\ast}(\widehat{\theta}_{t})|^{2}\right)^{\frac{1}{2}}\leq C(M,B,k,d)N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}.$
@@
Under the same condition as in Theorem 3.7, we have
$\left(\mathbb{E}|\widehat{\Pi}_{T}^{\ast}(\psi)-I|^{2}\right)^{\frac{1}{2}}\leq C(M,B,k,d,\epsilon)\bar{N}_{T}^{-1+\epsilon},\forall\epsilon>0.$
@@
(17) $\leq$ $C(M,B,k,d)\sum_{t=1}^{T}\frac{(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}}{\Omega_{T}}$
$\leq$ $C(M,B,k,d,\epsilon)\sum_{t=1}^{T}\frac{N_{t}^{\epsilon}}{\Omega_{T}}$
$\leq$ $C(M,B,k,d,\epsilon)(\frac{\Omega_{T}}{T})^{-1+\epsilon}=C(M,B,k,d,\epsilon)\bar{N}_{T}^{-1+\epsilon},$
@@ 3.4 Discrepancy between the MAMIS estimator and the auxiliary estimator
$\left(\mathbb{E}\left|\frac{1}{\Omega_{T}}\sum_{t=1}^{T}\sum_{i=1}^{N_{t}}\frac{\pi\left(X_{i}^{t}\right)\psi\left(X_{i}^{t}\right)}{\xi_{T}\left(X_{i}^{t}\right)}-\frac{1}{\Omega_{T}}\sum_{t=1}^{T}\sum_{i=1}^{N_{t}}\frac{\pi\left(X_{i}^{t}\right)\psi\left(X_{i}^{t}\right)}{q\left(X_{i}^{t},\theta^{*}\right)}\right|^{2}\right)^{\frac{1}{2}}$
@@
proceeds by decomposing $\bigg{|}\xi_{T}\left(X_{i}^{t}\right)-q\left(X_{i}^{t},\theta^{*}\right)\bigg{|}$ as
$\left|\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}q\left(X_{i}^{t},\widehat{\theta}_{l}\right)-q\left(X_{i}^{t},\theta^{\ast}\right)\right|$ $=\left|\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}\left[q\left(X_{i}^{t},\widehat{\theta}_{l}\right)-q\left(X_{i}^{t},\theta^{\ast}\right)\right]\right|$
@@
Consequently, from Eq. 20 and Theorem 3.3, we can bound Eq. 19 by
$L\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}\left[C(M,B,k,d,D,\eta)N_{l}^{-1}(\log N_{l})^{\frac{3d}{2}-1}\exp{\{M(8vd\log N_{l})^{\frac{k}{2}}\}}\right]$
$\leq$ $C(M,B,k,d,D,L,\eta,\epsilon)\sum_{l=1}^{T}\frac{N_{l}^{\epsilon}}{\Omega_{T}}$
$\leq$ $C(M,B,k,d,D,L,\eta,\epsilon)\bar{N}_{T}^{-1+\epsilon},\forall\epsilon>0,$
where we apply the Jensen’s inequality to the concave function $f(x)=x^{\epsilon}$ in the last inequality. As a result, we can also bound the right hand side of Eq. 18 by $C(M,B,k,d,D,L,\eta,\epsilon)\bar{N}_{T}^{-1+\epsilon}$.
@@
Under Section 3, we have
$\left(\mathbb{E}\left|\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-\widehat{\Pi}_{T}^{*}(\psi)\right|^{2}\right)^{\frac{1}{2}}\leq C(M,B,k,d,D,L,\eta,\epsilon)\Big{(}\bar{N}_{T}\Big{)}^{-1+\epsilon},\forall\epsilon>0.$
@@
$\leq C(M,B,k,d,D,L,\eta,\epsilon)\bar{N}_{T}^{-1+\epsilon}.$
@@ 4 Convergence Rate of the MC-based MAMIS Estimator
where $\widehat{I}_{n}(f)$ is given in Eq. 8 and the constant in big-$\mathcal{O}$ bound depends on $q,A$.
@@
By using this lemma and following an analysis process similar to that in the QMC case, we can obtain the convergence order result of the MC-based MAMIS estimator.
@@
$\Big{(}\mathbb{E}\Big{|}\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I\Big{|}^{2}\Big{)}^{\frac{1}{2}}\leq C(\eta,A)\bar{N}_{T}^{-1/2},$
where the constant $C(\eta,A)$ depends on $\eta,A$ and $\eta$ appears in Eq. 12.

### v2

@@ Abstract
This article investigates the integration of quasi-Monte Carlo (QMC) methods using Adaptive Multiple Importance Sampling (AMIS). Traditional Importance Sampling (IS) often suffers from poor performance since it heavily relies on the choice of the proposal distributions. The AMIS and the Modified version of AMIS (MAMIS) address this by iteratively refining proposal distributions and reusing all past samples through a recycling strategy. We introduce randomized QMC (RQMC) methods into the MAMIS, achieving higher convergence rates compared to Monte Carlo (MC) methods. Our main contributions include a detailed convergence analysis of the MAMIS estimator under the RQMC setting. Specifically, we establish the $L^{q}$ $(q\geq 2)$ error bound for the RQMC-based estimator using a smoothed projection method, which enables us to apply the Hölder’s inequality in the error analysis of the RQMC-based MAMIS estimator. As a result, we prove that the root mean square error of the RQMC-based MAMIS estimator converges at a rate of $\mathcal{O}(\bar{N}_{T}^{-1+\epsilon})$, where $\bar{N}_{T}$ is the average number of samples used in each step over $T$ iterations and $\epsilon>0$ is arbitrarily small. Numerical experiments validate the effectiveness of our method, including mixtures of Gaussians, a banana-shaped model, and Bayesian Logistic regression.
@@ 2.1 The modified version of AMIS
Utilizing the samples $X_{i}^{T}$ with weights $\omega_{i}^{T}$ obtained at the final stage $T$ in Algorithm 1, the AIS estimator is formulated as
@@
Marin et al. [15] also emphasized this recycling strategy, but instead of reweighting the samples at each stage $t$ to update the parameters, a simpler and more analytically tractable method was suggested. The resulting Modified version of AMIS algorithm is exposed in Algorithm 2. We use the assumption in [15], that is, under a criterion such as Kullback-Leibler divergence, variance, or others, the parameter $\theta^{\ast}$ which we desire $\widehat{\theta}_{t}$ to approximate, can be written as
@@ 3 Convergence rate of the RQMC-based MAMIS estimator
$\Big{(}\mathbb{E}\Big{|}\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I\Big{|}^{2}\Big{)}^{\frac{1}{2}}=\mathcal{O}(\bar{N}_{T}^{-1+\epsilon}),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@
$\left(\mathbb{E}|\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I|^{2}\right)^{\frac{1}{2}}=\mathcal{O}(N^{-1+\epsilon}),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@
Unfortunately, this corollary does not include the relationship between error and the number of iteration $T$. This is because $T$ actually appears in the constant term. In order to estimate the sum of errors at each stage, we use the same function class $G_{e}(M,B,k)$ to encompass all functions ${\pi\psi}/{q(\cdot,\theta^{\ast})}\circ T_{\theta},{\pi h_{j}}/{q(\cdot,\theta^{\ast})}\circ T_{\theta}$, which eliminated the influence of $T$ in the constant term. Moreover, Theorem 3.1 is not entirely devoid of the impact of $T$. When we use a strictly increasing sample size $\{N_{t}\}$, such as $N_{t}=Nt$, we have $\bar{N}_{T}=\frac{N(T+1)}{2}$. Applying Theorem 3.1 yields an error order as $\mathcal{O}((NT)^{-1+\epsilon})$. {Remark} The convergence order in Theorem 3.1 still holds for the case where the integrand $\psi$ is a finite-dimensional vector-valued function. This can be shown by applying Theorem 3.1 to each component function of $\psi$ and noting that the square of the vector norm is equal to the sum of the squares of its components. In this case, the constant term in Theorem 3.1 will also be related to the dimension of the vector-valued function.
@@ 3.1 Convergence of the parameters $\widehat{\theta}_{t}$
$\left(\mathbb{E}{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\right)^{\frac{1}{q}}=\mathcal{O}\left(N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(4qd\log N_{t})^{\frac{k}{2}}\}}\right),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,q$. Thus we can obtain
(14) $\left(\mathbb{E}{\lVert\widehat{\theta}_{t+1}-\theta^{\ast}\rVert}^{q}\right)^{\frac{1}{q}}=\mathcal{O}\left(N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(4qd\log N_{t})^{\frac{k}{2}}\}}\right),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,D,q$.
@@
From Eq. 13, the conditional expectation
$\mathbb{E}\left[{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\bigg{|}\widehat{\theta_{t}}\right]$
$=$ $\mathbb{E}\left[\Big{|}\frac{1}{N_{t}}\sum_{i=1}^{N_{t}}h_{\widehat{\theta}_{t}}^{(j)}\circ F_{\widehat{\theta}_{t}}(u_{i}^{t})-\int_{[0,1)^{d}}h_{\widehat{\theta}_{t}}^{(j)}\circ F_{\widehat{\theta}_{t}}(\bm{u})\mathrm{d}\bm{u}\Big{|}^{q}\bigg{|}\widehat{\theta}_{t}\right]$
@@
$\lesssim$ $N_{t}^{-q}(\log N_{t})^{q(\frac{3d}{2}-1)}\exp{\{qM(4qd\log N_{t})^{\frac{k}{2}}\}}$
for any $\theta\in\Theta$, where the symbol $\lesssim$ is used for hiding a constant independently of $N_{t}$ and Eq. 15 employs the independence between the RQMC point set $\{u_{i}^{t}\}_{i=1}^{N_{t}}$ and $\widehat{\theta}_{t}$. Thus we obtain
$\mathbb{E}\left[{\left|\widehat{\theta}_{t+1,j}-\theta^{\ast}_{j}\right|}^{q}\Big{|}\widehat{\theta_{t}}\right]\lesssim N_{t}^{-q}(\log N_{t})^{q(\frac{3d}{2}-1)}\exp{\{qM(4qd\log N_{t})^{\frac{k}{2}}\}}.\$
@@ 3.2 On the property of operator $I_{\psi}^{\ast}$
$\Big{(}\mathbb{E}\big{|}I_{\psi}^{\ast}(\widehat{\theta}_{t})-I_{\psi}^{\ast}(\theta^{\ast})\big{|}^{2}\Big{)}^{\frac{1}{2}}=\mathcal{O}\left(N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}\right),$
where the constant in the big-$\mathcal{O}$ constant depends on $M,B,k,d,D,L$.
@@ 3.3 Convergence of the auxiliary estimator
$\left(\mathbb{E}|\widehat{\pi}_{t}^{\ast}(\psi)-I_{\psi}^{\ast}(\widehat{\theta}_{t})|^{2}\right)^{\frac{1}{2}}=\mathcal{O}\left(N_{t}^{-1}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}\right),$
where the constant in the big-$\mathcal{O}$ constant depends on $M,B,k,d$.
@@
Under the same condition as in Theorem 3.7, for any $\epsilon>0$, we have
$\left(\mathbb{E}|\widehat{\Pi}_{T}^{\ast}(\psi)-I|^{2}\right)^{\frac{1}{2}}=\mathcal{O}\left(\bar{N}_{T}^{-1+\epsilon}\right),$
where the constant in the big-$\mathcal{O}$ constant depends on $M,B,k,d,\epsilon$.
@@
(17) $\lesssim$ $\frac{1}{{\Omega_{T}}}\sum_{t=1}^{T}(\log N_{t})^{\frac{3d}{2}-1}\exp{\{M(8d\log N_{t})^{\frac{k}{2}}\}}$
$\lesssim$ $\frac{1}{{\Omega_{T}}}\sum_{t=1}^{T}N_{t}^{\epsilon}\lesssim(\frac{\Omega_{T}}{T})^{-1+\epsilon}=\bar{N}_{T}^{-1+\epsilon},$
@@ 3.4 Discrepancy between the MAMIS estimator and the auxiliary estimator
$\left(\mathbb{E}\left|\widehat{\Pi}_{T}^{\text{MAMIS}}(\psi)-\widehat{\Pi}_{T}^{*}(\psi)\right|^{2}\right)^{\frac{1}{2}}$
$=$ $\left(\mathbb{E}\left|\frac{1}{\Omega_{T}}\sum_{t=1}^{T}\sum_{i=1}^{N_{t}}\frac{\pi\left(X_{i}^{t}\right)\psi\left(X_{i}^{t}\right)}{\xi_{T}\left(X_{i}^{t}\right)}-\frac{1}{\Omega_{T}}\sum_{t=1}^{T}\sum_{i=1}^{N_{t}}\frac{\pi\left(X_{i}^{t}\right)\psi\left(X_{i}^{t}\right)}{q\left(X_{i}^{t},\theta^{*}\right)}\right|^{2}\right)^{\frac{1}{2}}$
@@
proceeds by
$\left|\xi_{T}\left(X_{i}^{t}\right)-q\left(X_{i}^{t},\theta^{*}\right)\right|$ $=\left|\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}q\left(X_{i}^{t},\widehat{\theta}_{l}\right)-q\left(X_{i}^{t},\theta^{\ast}\right)\right|$
$=\left|\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}\left[q\left(X_{i}^{t},\widehat{\theta}_{l}\right)-q\left(X_{i}^{t},\theta^{\ast}\right)\right]\right|$
@@
Consequently, from Eq. 20 and Theorem 3.3, we can bound Eq. 19 as follows.
$\left({\mathbb{E}\left|\frac{g(X_{i}^{t})}{\xi_{T}\left(X_{i}^{t}\right)}\left(\xi_{T}\left(X_{i}^{t}\right)-q\left(X_{i}^{t},\theta^{\ast}\right)\right)\right|^{2}}\right)^{\frac{1}{2}}$
$\lesssim$ $\sum_{l=1}^{T}\frac{N_{l}}{\Omega_{T}}\left[N_{l}^{-1}(\log N_{l})^{\frac{3d}{2}-1}\exp{\{M(8vd\log N_{l})^{\frac{k}{2}}\}}\right]$
$\lesssim$ $\frac{1}{\Omega_{T}}\sum_{l=1}^{T}N_{l}^{\epsilon}\lesssim(\frac{\Omega_{T}}{T})^{-1+\epsilon}=\bar{N}_{T}^{-1+\epsilon},\forall\epsilon>0,$
where we apply the Jensen’s inequality to the concave function $f(x)=x^{\epsilon}$ in the last inequality. As a result, we can also bound the right hand side of Eq. 18 by $\mathcal{O}(\bar{N}_{T}^{-1+\epsilon})$.
@@
Under Section 3, for any $\epsilon>0$, we have
$\left(\mathbb{E}\left|\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-\widehat{\Pi}_{T}^{*}(\psi)\right|^{2}\right)^{\frac{1}{2}}=\mathcal{O}(\bar{N}_{T}^{-1+\epsilon}),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@
$=\mathcal{O}(\bar{N}_{T}^{-1+\epsilon}),$
where the constant in the big-$\mathcal{O}$ bound depends on $M,B,k,d,D,L,\eta,\epsilon$.
@@ 4 Convergence Rate of the MC-based MAMIS Estimator
where $\widehat{I}_{n}(f)$ is given in Eq. 8 and the constant in the big-$\mathcal{O}$ bound depends on $q,A$.
@@
By using this lemma and following an analysis process similar to that in the RQMC case, we can obtain the convergence order result of the MC-based MAMIS estimator.
@@
$\Big{(}\mathbb{E}\Big{|}\widehat{\Pi}_{T}^{\mathrm{MAMIS}}(\psi)-I\Big{|}^{2}\Big{)}^{\frac{1}{2}}=\mathcal{O}(\bar{N}_{T}^{-1/2}),$
where the constant in the big-$\mathcal{O}$ bound depends on $L,\eta,A$.